# ProjectTest: A Project-level LLM Unit Test Generation Benchmark and Impact of Error Fixing Mechanisms

### v3

@@ 1 Introduction
We evaluate nine frontier LLMs, such as Claude-3.5-Sonnet Anthropic (2024), Gemini-1.5-Pro Team et al. (2024b), and GPT-o1, on ProjectTest and conduct comprehensive error analyses. We find that all tested frontier LLMs perform moderately on ProjectTest on Python and Java, highlighting the difficulty of ProjectTest. We also observe that different LLMs have different language-level expertise. Claude-3.4-Sonnet ranks first in Java, while GPT-o1 ranks first in JavaScript. Among three programming languages, Java is the most difficult language, primarily due to stricter syntax. Among all the tested models, GPT-o1 performs the best in general, especially in JavaScript.
@@ 4.1 Models
We evaluate five close-sourced models: GPT-o1, Gemini-1.5-Pro Team et al. (2024b), Claude-3.5-Sonnet-20241022 (Claude-3.5-Sonnet) Anthropic (2024), GPT-4-Turbo Achiam et al. (2023) and GPT-3.5-Turbo, and four open-sourced models: CodeQwen1.5-7B-Chat (CodeQwen1.5) Bai et al. (2023), DeepSeek-Coder-6.7b-Instruct (DeepSeek-Coder) Guo et al. (2024); Zhu et al. (2024), CodeLlama-7b-Instruct-hf (CodeLlama) Roziere et al. (2023), and CodeGemma-7b-it (CodeGemma) Team et al. (2024a). Detailed information is in Appendix B.2.
@@ 5.1 Main Results
Among three programming languages, Java is the most difficult language, primarily due to stricter syntax. Many models fail to generate valid Java code, leading to low compilation rates and execution coverage. Among all the evaluated models, GPT-o1 performs the best in general, especially in JavaScript. CodeLlama and CodeGemma have the worst general performance. We also observe that some models tend to generate more unit tests. However, generating more unit tests does not necessarily lead to better coverage rates. For example, Gemini-1.5-Pro tends to generate the most unit tests but does not obtain the best coverage rate. Additionally, we observe that sometimes the open-source model can even outperform some closed-source models. For example, DeepSeek-Coder works better than GPT-3.5-Turbo on Python and JavaScript. Finally, we confirmed from such results that dependencies exist in metrics. On Java, models like CodeQwen1.5, CodeLlama, and CodeGemma fail to generate compilable unit tests, resulting in the lowest correctness rates and coverage rates.
@@ 5.2 Manual Fixing Results
Among all programming languages, Java benefits the most from manual fixing. In the case of vanilla unit tests, Java exhibits the lowest compilation rates, making it particularly challenging. However, after manual fixing, Java shows the most substantial improvement, highlighting the potential of LLMs for Java after fixing compilation and cascade errors. Among all models, GPT-o1 still performs the best after manual fixing, and CodeLlama and CodeGemma still exhibit the worst general performance. Gemini-1.5-Pro shows the best coverage improvement overall, indicating its strong potential for better unit test generation once compilation and cascade errors are corrected. Finally, we observe that certain models that initially underperform can match or surpass stronger models after manual fixing. For example, in Java, CodeQwen1.5 outperforms DeepSeek-Coder and is now on par with GPT-4-Turbo. In Python, Gemini-1.5-Pro surpasses CodeQwen1.5-7B-Chat, showing better potential after manual fixing. On JavaScript, GPT-3.5-Turbo has reached parity with DeepSeek-Coder.
@@ B.2 Models
Close-sourced Gemini-1.5-Pro - https://ai.google.dev/gemini-api/docs/models/gemini#gemini-1.5-pro

### v4

@@ 1 Introduction
We evaluate nine frontier LLMs, such as Claude-3.5-Sonnet Anthropic (2024), Gemini-2.0-Flash Team et al. (2024b), and GPT-o1, on ProjectTest and conduct comprehensive error analyses. We find that all tested frontier LLMs perform moderately on ProjectTest on Python and Java, highlighting the difficulty of ProjectTest. We also observe that different LLMs have different language-level expertise. Claude-3.4-Sonnet ranks first in Java, while GPT-o1 ranks first in JavaScript. Among three programming languages, Java is the most difficult language, primarily due to stricter syntax. Among all the tested models, GPT-o1 performs the best in general, especially in JavaScript.
@@ 4.1 Models
We evaluate five close-sourced models: GPT-o1, Gemini-2.0-Flash-Exp Team et al. (2024b), Claude-3.5-Sonnet-20241022 (Claude-3.5-Sonnet) Anthropic (2024), GPT-4-Turbo Achiam et al. (2023) and GPT-3.5-Turbo, and four open-sourced models: CodeQwen1.5-7B-Chat (CodeQwen1.5) Bai et al. (2023), DeepSeek-Coder-6.7b-Instruct (DeepSeek-Coder) Guo et al. (2024); Zhu et al. (2024), CodeLlama-7b-Instruct-hf (CodeLlama) Roziere et al. (2023), and CodeGemma-7b-it (CodeGemma) Team et al. (2024a). Detailed information is in Appendix B.2.
@@ 5.1 Main Results
Among three programming languages, Java is the most difficult language, primarily due to stricter syntax. Many models fail to generate valid Java code, leading to low compilation rates and execution coverage. Among all the evaluated models, GPT-o1 performs the best in general, especially in JavaScript. CodeLlama and CodeGemma have the worst general performance. We also observe that some models tend to generate more unit tests. However, generating more unit tests does not necessarily lead to better coverage rates. For example, Gemini-2.0-Flash tends to generate the most unit tests but does not obtain the best coverage rate. Additionally, we observe that sometimes the open-source model can even outperform some closed-source models. For example, DeepSeek-Coder works better than GPT-3.5-Turbo on Python and JavaScript. Finally, we confirmed from such results that dependencies exist in metrics. On Java, models like CodeQwen1.5, CodeLlama, and CodeGemma fail to generate compilable unit tests, resulting in the lowest correctness rates and coverage rates.
@@ 5.2 Manual Fixing Results
Among all programming languages, Java benefits the most from manual fixing. In the case of vanilla unit tests, Java exhibits the lowest compilation rates, making it particularly challenging. However, after manual fixing, Java shows the most substantial improvement, highlighting the potential of LLMs for Java after fixing compilation and cascade errors. Among all models, GPT-o1 still performs the best after manual fixing, and CodeLlama and CodeGemma still exhibit the worst general performance. Gemini-2.0-Flash shows the best coverage improvement overall, indicating its strong potential for better unit test generation once compilation and cascade errors are corrected. Finally, we observe that certain models that initially underperform can match or surpass stronger models after manual fixing. For example, in Java, CodeQwen1.5 outperforms DeepSeek-Coder and is now on par with GPT-4-Turbo. In Python, Gemini-2.0-Flash surpasses CodeQwen1.5-7B-Chat, showing better potential after manual fixing. On JavaScript, GPT-3.5-Turbo has reached parity with DeepSeek-Coder.
@@ B.2 Models
Close-sourced Gemini-2.0-Flash - https://ai.google.dev/gemini-api/docs/models/gemini#gemini-2.0-Flash